OpenStudy (anonymous):

How do you factor x(x+3)+2(x+3)?

OpenStudy (anonymous):

x(x+3)+2(x+3) = x²+3x + 2x + 6 = x²+5x+6=(x+2)(x+3)

OpenStudy (urimare):

(x+3)(x+2)

OpenStudy (anonymous):

Factor by grouping: Treat the problem as 2 terms notice that both terms have x+3 as factors factor out x+3 from the first term you will have an x left over from your 2nd term you will have a 2 left over (x+3) ( x + 2 )